# Creation of Wavepackets for Quantum Chromodynamics on Quantum Computers

### v1

@@ V.1 Staggered Fermions in Three Dimensions
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}\\ +\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}d(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}u(\mathbf{r})+iu^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}d(\mathbf{r})-id^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}u(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]},\end{gathered}$ (81)
@@
$\Delta_{\hat{\mathbf{x}}}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{x}})-q(\mathbf{r}-\hat{\mathbf{x}})\big{]},$ $\qquad\Delta_{\hat{\mathbf{x}}}^{2}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{x}})+q(\mathbf{r}-\hat{\mathbf{x}})-2q(\mathbf{r})\big{]},$ (82)
$\Delta_{\hat{\mathbf{y}}}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{y}})-q(\mathbf{r}-\hat{\mathbf{y}})\big{]},$ $\qquad\Delta_{\hat{\mathbf{y}}}^{2}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{y}})+q(\mathbf{r}-\hat{\mathbf{y}})-2q(\mathbf{r})\big{]},$
@@
The main difference between the two identifications is that in the second one there are mixing terms between the two flavors already in the free theory. For this reason, in the following we consider only the first identification, equations (70) and (74).
@@ V.2 Interpolators in the Kogut-Susskind Formulation
$H_{\text{K}}$ $=\frac{1}{2a}\sum_{\mathbf{r},\hat{\mathbf{n}}}\big{[}\xi^{\dagger}(\mathbf{r})U(\mathbf{r},\hat{\mathbf{n}})\xi(\mathbf{r}+\hat{\mathbf{n}})+\xi^{\dagger}(\mathbf{r}+\hat{\mathbf{n}})U^{\dagger}(\mathbf{r},\hat{\mathbf{n}})\xi(\mathbf{r})\big{]}\eta(\mathbf{r},\hat{\mathbf{n}}),$ (85)
@@
where $P_{\square}$ is a plaquette operator.
Our goal here is to find interpolating operators for the particles appearing in QCD, like the pion and the nucleons. Knowing the correspondence between the fields $u$ and $d$ and $\xi$, equations (70) and (74), we can use results from the literature on traditional lattice calculations [38, 39, 40, 41]. However, the symmetries of the formulation used in [38, 39, 40, 41] are quite different from the symmetries of the Kogut-Susskind Hamiltonian. In particular, cubic and flavor symmetries are intertwined between each other for staggered fermions, which implies that spin and isospin mix together. On the other hand, spin and isospin are among the main symmetries used to classify and identify particles. As a consequence, identification of particles in the Kogut-Susskind formulation is possible only near the continuum limit, and up to $O(a)$ errors.
As examples, consider the pions and the nucleons. Pion interpolators are
@@
where $C=-i\alpha_{y}$ is the charge-conjugation matrix in the Dirac representation. The neutron interpolator is obtained by exchanging $u$ and $d$.
These operators are not gauge invariant in the Kogut-Susskind formulation, because they involve products between $\xi$ operators at different lattice sites as a consequence of the identifications in equation (70). We need to insert appropriate link operators to make them gauge invariant. The mesonic operators contain terms like $\xi^{\dagger}(2\mathbf{r})\xi(2\mathbf{r}+\hat{\mathbf{x}})$, which is made gauge invariant by $\xi^{\dagger}(2\mathbf{r})U(2\mathbf{r},\hat{\mathbf{x}})\xi(2\mathbf{r}+\hat{\mathbf{x}})$. The baryonic operators contain terms like $\epsilon_{abc}\xi_{a}(2\mathbf{r})\xi_{b}(2\mathbf{r}+\hat{\mathbf{x}})\xi_{c}(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}})$, which is made gauge invariant by
$\epsilon_{abc}U_{aa^{\prime}}(2\mathbf{r}+\hat{\mathbf{x}},-\hat{\mathbf{x}})\xi_{a^{\prime}}(2\mathbf{r})\xi_{b}(2\mathbf{r}+\hat{\mathbf{x}})U_{cc^{\prime}}(2\mathbf{r}+\hat{\mathbf{x}},\hat{\mathbf{y}})U_{c^{\prime}c^{\prime\prime}}(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}},\hat{\mathbf{x}})\xi_{c^{\prime}}(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}).$ (90)
Following the examples given in the previous section on one-dimensional lattices, we can map the interpolating operators into qubit operators, and decompose them using LCU. The link operators $U_{\alpha\beta}$ take the same form as in the one dimensional case. The main difference with the previous section is in the encoding of the fermionic field $\xi$. According to [42], it is computationally more convenient to use the generalized superfast encoding (GSE) instead of a Jordan-Wigner transformation in three dimensions. We review the GSE in the appendix, where we also explain how to encode odd fermionic operators that are necessary to implement the nucleon interpolators.
We conclude this section with some considerations on the success probability $\rho$ in equation (7). The factor $C$ depends on the number of terms in the interpolating operator and on its mass dimension. Each quark field contains a factor $1/a^{1.5}$ in the continuum normalization. Counting also the numerical coefficients in the expansions of the interpolators in terms of $\xi$, we obtain approximately $C_{\pi^{\pm}}\sim 432/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}\sim 11500/a^{4.5}$. The large numerical coefficients appearing here are a consequence of choosing Kogut-Susskind fermions. For a comparison, the corresponding coefficients using Wilson fermions are $C_{\pi^{\pm}}=12/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}=24/a^{4.5}$.
@@ VI Conclusions
In reference [25] we showed how to create wavepackets from the vacuum in scalar field theories. Here we focused on lattice gauge theories, and increased gradually the level of difficulty, starting from a $U(1)$ theory in one dimension and ending with QCD in three dimensions with two flavors of quarks. We considered different kind of particles, namely pions and nucleons in QCD.
@@
Three ingredients are necessary to apply our method. First, we need to prepare the vacuum of the theory. This can be done with variational methods, like [43, 24], or with an adiabatic transformation [44]. Second, we need to know the spectrum of the theory. Near the continuum limit, this amounts to know the particle content and the particle masses. Finally, we need interpolating operators for the particles we aim at. These operators are built on a symmetry basis, and they should carry the quantum numbers of the particles. Our work is further extensible to other theories, formulations and particles, with the only requirement that there has to be a mass gap.
@@
The creation of a wavepacket with our strategy requires the use of LCU. This technique succeeds with a probability $\rho$ that depends on the theory, on the wavepacket, and on the particle. If we want to prepare $n$ wavepackets, the initial state preparations needs to be repeated on average $O(1/\rho^{n})$ times to obtain the correct state for the simulation, so we can create only a small number of incoming wavepackets. Fortunately, $n=2$ is the most relevant scenario. The success probability seems to be the main limitation of our approach, but at least it does not affect the circuit depth, which is a favorable feature.
@@
We chose to work in the Kogut-Susskind framework. For QCD, we imported interpolators for the pions and the nucleons from the literature on traditional lattice calculations. We transformed these operators into operators in the Kogut-Susskind formulation, but this rises some problems. There are many possible choices for these transformations, with differences of the order of the lattice spacing, and none of them respects the symmetries that determine the particles in the continuum. As a consequence, particle identification is more problematic, a well-known downside of staggered fermions. Furthermore, the success probability suffers significantly from the transformation of staggered fermions into Dirac spinors. The interpolating operators in the Kogut-Susskind formulation become long linear expansions of operators, with insertion of gauge fields to achieve gauge invariance. The resulting overheads in the success probability are $\sim 10^{4}$ for the pions, and $\sim 10^{8}$ for the nucleons. If we use Wilson fermions, the equivalent overheads are $\sim 10^{2}$.
@@
In the future, we aim to provide more precise estimates on the resources needed, especially in the cases that have already been tested [23, 24] to make a comparison between the different strategies. If the hardware is mature enough, it would be useful to test the method on a real quantum computer. It would also be important to have asymptotical estimates of the other approaches in the literature to determine which one is the most promising.
@@ VII Acknowledgments
We thank the support from FCT – Fundação para a Ciência e a Tecnologia (Portugal), namely through project UIDB/04540/2020, as well as from projects QuantHEP and HQCC supported by the EU QuantERA ERA-NET Cofund in Quantum Technologies and by FCT (QuantERA/004/2021). Finally, Matteo Turco thanks FCT for support through the grant PRT/BD/154668/2022.
Appendix A Generalised Superfast Encoding of Fermions
We review the generalised superfast encoding (GSE) of fermions into qubits [45]. In [42], this encoding was used to estimate the resource requirements for quantum simulation of lattice gauge theories such as quantum chromodynamics, showing a substantial improvement with respect to a Jordan-Wigner transformation. As presented in [45], the GSE allows for the representation of even fermionic operators only. As odd operators are necessary for the preparation of fermionic-particle wavepackets, here we also show how they can be represented in the GSE. In this appendix we drop the Einstein convention for repeated indices.
@@ A.2 Even Operators in the GSE
=$\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ forall loops }\zeta\}.$ (108)
@@ A.3 Odd Operators in the GSE
Consider again a graph as in the previous section, with every vertex of even degree, and take two vertices $S$ and $T$ connected by an edge. For every vertex $I$ in the graph take a path $\zeta_{SI}$ from $S$ to $I$ that does not use the edge $(S,T)$, with $\zeta_{SI}(1)=S$ and $\zeta_{SI}(s)=I$. We can always find such paths following the Eulerian cycle $\eta$. Consider now the local Majorana mode on $S$ attached to the edge $(S,T)$, which we can call $\gamma_{S1}$ (up to relabelling of the modes). Motivated by
@@
However, on the subspace =$\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\}$, we have
@@
+=$\mathcal{L}^{+}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=\tilde{A}(\eta)$ (124)
@@
=$\mathcal{L}^{-}=\{|\psi\rangle:\bar{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=-\bar{A}(\eta).$ (127)
@@ A.4 GSE for Hypercubic Lattices
In this work we use the GSE for $SU(N)$ gauge theories on hypercubic lattices with one flavour of fermions, which means that we have $N$ independent fermionic modes on each site. In the one-dimensional case, the GSE essentially reduces to the Jordan-Wigner transformation. In $d$ dimensions we use $N+d-1$ qubits per lattice site. As an example, figure 4 depicts two neighbouring sites in an $SU(3)$ theory on a square lattice. The gray lines represent physical interaction terms in the Hamiltonian, and there is no such line between modes on the same site. The thick lines are edges of the graph we use for our encoding, and we also show the associated local Majorana modes on each mode. To obtain an interaction term between two modes, we simply choose the shortest path on the graph connecting the two modes, and the resulting encoded operator is in general a string of operators acting on the qubits involved in the path. Two out of three fermionic modes use one qubit each, the other modes use two qubits each. If mode $I$ uses one qubit, we take as local Majorana modes

### v2

@@ V.1 Staggered fermions in three dimensions
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}a^{3}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}\\ +\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}d(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}u(\mathbf{r})+iu^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}d(\mathbf{r})-id^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}u(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]},\end{gathered}$ (83)
@@
[…]
@@
The main difference between the two identifications is that the first one is slightly more involved, but it avoids terms that mix the two flavours in the free theory. The second one is simpler, but the identified flavours are mixed. We give them both because it is not clear which one is more convenient. However, in the following, we consider only the first identification, equations (72) and (76), since it allows for a better identification of the flavours.
@@ V.2 Interpolators in the Kogut-Susskind formulation
[…]
@@
where $P_{\square}$ is a plaquette operator. The fields $\xi$, $U$ and $E$ are the $SU(3)$ fields discussed in the previous section. The encoding of the link Hilbert spaces does not differ from the one-dimensional case, while for the fermions it is more convenient to use a local encoding such as the generalized superfast encoding [75, 76]. This encoding was given only for operators made of products of an even number of fermionic operators. To create wave packets of a fermionic particle, such as the nucleons, we need to encode also odd fermionic operators. In the appendix we review the generalized superfast encoding and extend it to odd operators. In practice, the encoded versions of the operators $\xi$ and $\xi^{\dagger}$ are made of operators $\sigma^{\pm}$ attached to a string of Pauli operators $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$ on other qubits. This is qualitatively similar to a Jordan-Wigner transformation, except that the strings of Pauli operators have different shapes depending on the details of the mapping.
Our goal here is to show how to write interpolating operators for the particles appearing in QCD, like the pion and the nucleons, in terms of qubit operators. Knowing the correspondence between the fields $u$ and $d$ and $\xi$, equations (72) and (76), we can use results from the literature on traditional lattice calculations [77, 78, 79, 80]. However, the symmetries of the formulation used in [77, 78, 79, 80] are different from the symmetries of the Kogut-Susskind Hamiltonian. In particular, cubic and flavor symmetries are intertwined between each other for staggered fermions, which implies that spin and isospin mix together. On the other hand, spin and isospin are among the main symmetries used to classify and identify particles. As a consequence, identification of particles in the Kogut-Susskind formulation is possible only near the continuum limit, and up to $O(a)$ errors.
Pion interpolators are
@@
where $\mathcal{C}=-i\alpha_{y}$ is the charge-conjugation matrix in the Dirac representation. The neutron interpolator is obtained by exchanging $u$ and $d$.
[…]
The necessity of inserting link operators to restore gauge invariance, together with the transformations (72) and (76), makes writing interpolators in the Kogut-Susskind formulation extremely cumbersome. This becomes particularly relevant when we calculate the quantities $C$ for these operators: We obtain approximately $C_{\pi^{\pm}}\sim 432/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}\sim 11500/a^{4.5}$. In a formulation such as one with Wilson fermions, in which each site contains the two spinors $u$ and $d$, the corresponding quantities would be $C_{\pi^{\pm}}=12/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}=24/a^{4.5}$.
@@ VI Conclusions
Quantum computation opens the possibility to study real-time evolution of complex many-body quantum systems in a general and systematic way. However, much work is required on the experimental and theoretical sides to achieve meaningful simulations that are inaccessible with traditional techniques. In this manuscript, we aimed at solving the problem of preparing initial states suitable for scattering simulation, consisting of two wave packets moving on top of the vacuum. Simulating scattering of wave packets not only allows for the computation of physical quantities to be compared with experimental results, but also for disclosing what happens inside the collision.
@@
Our method is a practical application of the Haag-Ruelle scattering theory developed in the context of axiomatic quantum field theory more than half a century ago. In reference [56] we showed how to create wave packets from the vacuum in scalar-field theories. Here we focused on lattice gauge theories, and increased gradually the level of difficulty, starting from a $U(1)$ theory in one dimension and ending with QCD in three dimensions with two flavors of quarks. We considered different kinds of particles, namely pions and nucleons in QCD.
@@
Three ingredients are necessary to apply our method. First, we need to prepare the vacuum of the theory. This could be done with variational methods, like [81, 52], or with an adiabatic transformation [82]. Second, we need to know the spectrum of the theory. Near the continuum limit, this amounts to know the particle content and the particle masses. Finally, we need interpolating operators for the particles we aim at. These operators are built on a symmetry basis, and they should carry the quantum numbers of the particles. Our work is further extensible to other theories, formulations and particles, with the only requirement that there has to be a mass gap.
@@
The creation of a wave packets with our strategy requires the use of LCU. This technique succeeds with a probability $\rho$ that depends on the theory, on the wave packet, and on the particle. If we want to prepare $n$ wave packets, the initial state preparations needs to be repeated on average $O(1/\rho^{n})$ times to obtain the correct state for the simulation, so we can create only a small number of incoming wave packets. Fortunately, $n=2$ is the most relevant case. The success probability seems to be the main limitation of our approach, but at least it does not affect the circuit depth, which is a favorable feature.
@@
We chose to work in the Kogut-Susskind framework with staggered fermions. For QCD, we imported interpolators for the pions and the nucleons from the literature on traditional lattice calculations. We transformed these operators into operators in the staggered-fermion formulation, but this raises some problems. There are many possible choices for these transformations, with differences of the order of the lattice spacing, and none of them respects the symmetries that determine the particles in the continuum. As a consequence, particle identification is more problematic, a well-known downside of staggered fermions. Furthermore, the success probability suffers significantly from the transformation of staggered fermions into Dirac spinors. The interpolating operators in the staggered-fermion formulation become long linear expansions of operators, with insertion of gauge fields to achieve gauge invariance. The resulting overheads in the success probability are $\sim 10^{4}$ for the pions, and $\sim 10^{8}$ for the nucleons. If we use Wilson fermions, the equivalent overheads are $\sim 10^{2}$.
@@
In the future, the main issue that should be addressed is the success probability. Finding a different method to implement the creation operators with good probability would significantly improve our approach. Furthermore, it would be interesting to have detailed resource estimates, especially in the cases that have already been tested [50, 52, 51] to make a comparison between the different strategies. When the hardware is mature enough, it will be useful to test the method on a real quantum computer. It would also be important to have asymptotic estimates of the other approaches in the literature to determine which one is the most promising.
@@ VII Acknowledgments
We are grateful for the support from FCT – Fundação para a Ciência e a Tecnologia (Portugal) – namely through Project No. UIDB/04540/2020 and Contract LA/P/0095/ 2020, as well as from projects QuantHEP and HQCC supported by the EU QuantERA ERA-NET Cofund in Quantum Technologies and by FCT (QuantERA/004/2021). Finally, M. T. thanks FCT for support through Grant No. PRT/BD/154668/2022. G. Q. acknowledges the support from FCT, through projects CEECIND/ 02474/2018, 2024.04456.CERN and FCT-Mobility/ 1312232346/2024-25.
Appendix A generalized superfast encoding of fermions
We review the generalized superfast encoding (GSE) of fermions into qubits [75]. In [76], this encoding was used to estimate the resource requirements for quantum simulation of lattice gauge theories such as quantum chromodynamics, showing a substantial improvement with respect to a Jordan-Wigner transformation. As presented in [75], the GSE allows for the representation of even fermionic operators only. As odd operators are necessary for the preparation of fermionic-particle wave packets, here we also show how they can be represented in the GSE. In this appendix we drop the Einstein convention for repeated indices.
@@ A.2 Even operators in the GSE
$\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ for all loops }\zeta\}.$ (110)
@@ A.3 Odd operators in the GSE
Consider again a graph as in the previous section, with every vertex of even degree, and take two vertices $S$ and $T$ connected by an edge. For every vertex $I$ in the graph take a path $\zeta_{SI}$ from $S$ to $I$ that does not use the edge $(S,T)$, with $\zeta_{SI}(1)=S$ and $\zeta_{SI}(s)=I$. We can always find such paths following the Eulerian cycle $\eta$. Consider now the local Majorana mode on $S$ attached to the edge $(S,T)$, which we can call $\gamma_{S1}$ (up to relabeling of the modes). Motivated by
@@
However, on the subspace $\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\}$, we have
@@
$\mathcal{L}^{+}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=\tilde{A}(\eta)$ (126)
@@
$\mathcal{L}^{-}=\{|\psi\rangle:\bar{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=-\bar{A}(\eta).$ (129)
@@ A.4 GSE for hypercubic lattices
In this work we use the GSE for $SU(N)$ gauge theories on hypercubic lattices with one flavour of fermions, which means that we have $N$ independent fermionic modes on each site. In the one-dimensional case, the GSE essentially reduces to the Jordan-Wigner transformation. In $d$ dimensions we use N+d1N+d-1 qubits per lattice site. As an example, figure 4 depicts two neighbouring sites in an $SU(3)$ theory on a square lattice. The gray lines represent physical interaction terms in the Hamiltonian, and there is no such line between modes on the same site. The thick lines are edges of the graph we use for our encoding, and we also show the associated local Majorana modes on each mode. To obtain an interaction term between two modes, we simply choose the shortest path on the graph connecting the two modes, and the resulting encoded operator is in general a string of operators acting on the qubits involved in the path. Two out of three fermionic modes use one qubit each, the other modes use two qubits each. If mode $I$ uses one qubit, we take as local Majorana modes